# Standing-Wave Dynamics in Low-Frequency Breathing of a Turbulent Separation Bubble

### v2

@@ 3.5 Standing Wave Model
In the experimental setup described in Chapter 2.1 the side walls are solid and not periodic. As with the periodic setup, the wavelength is constrained by the channel width, $L$, and the wavenumbers are discrete. The presence of a solid wall, however, introduces an additional constraint. In solid mechanics and acoustics, wave reflection at boundaries is a well-known phenomenon. A simple demonstration is the classic experiment where a wave pulse traveling along a string fixed at one end reflects upon reaching the boundary. Similarly, water waves hitting solid surfaces exhibit reflection, and in acoustics, echoes in large spaces provide an intuitive example of wave reflection. In compressible numerical simulations, boundary conditions incorporating reflection and refraction properties are standard. However, while our setup is incompressible, we still assume reflection of hydrodynamic waves at the walls.
@@
To derive the standing-wave model, we consider a wave with an arbitrary wavenumber $\beta$ that travels in positive spanwise direction and is reflected at the wall at $z=L$, undergoes a change in orientation that changes the sign of the wavenumber, and propagates to $z=0$, where it is reflected again. In the presence of reflecting boundaries at both ends, a wave is always accompanied by a reflected wave with the same wavelength but inverted orientation. The superposition of two waves of opposite directions with the same wavelength results in the formation of a standing wave.
The ansatz for the standing-wave model is then given as:
@@
that is constructed of two resolvent modes with opposite spanwise phase velocities. The same approach could be applied to the eigenmodes.
To incorporate spanwise boundary conditions, we enforce slip walls. The slip-wall boundary condition constrains the wall-normal velocity perturbation, $w^{\prime}_{sw}$, to zero at $z=0$ and $z=L$. This holds for all spanwise wavenumbers $\beta$ at $z=0$, but only for $\beta=n\pi/L$ , $n\in\mathbb{Z}$, at $z=L$. The allowed wavenumbers in the standing wave model are thus
@@
In summary, boundary conditions impose fundamental restrictions on the allowable wavelengths. Periodic sidewalls, often employed in numerical studies, permit only integer wavenumbers, with the maximum wavelength constrained by the channel width. By drawing an analogy to acoustics and water waves, we assume the sidewalls exhibit a reflective property. In this framework, resolvent-based waves undergo reflection at the sidewalls, leading to the formation of standing waves. The validity of this model and its agreement with experimental data will be demonstrated in the following chapter.
@@ 4.1 Empirical description of the TSB low-frequency dynamics
The SPOD spectrum of the streamwise velocity fluctuations measured by PIV on the symmetry plane is presented in Figure 6(a) which shows the mode energies of all SPOD modes $\lambda_{i}$. Figure 7(a) shows the energy share of each mode at every frequency which is calculated by dividing the mode energy by the sum of all modes at the given frequency. It reveals a leading mode that has a significantly higher energy share compared to the other modes. This is also known as high-gain separation and describes low-rank dynamics, meaning that the dynamics of the flow are dominated by one or very few coherent structures in the corresponding frequency range (SchmidtColonius2020). This mode concentrates the majority of its energy within the low-frequency range, specifically below $St=0.05$. The lowest frequency resolvable from our experimental data is $St=0.01$. A similar low-rank behaviour for low frequencies has been observed by WeissAIAAJ2022 using SPOD of wall-shear stress signals. Instead of a single dominant characteristic frequency, the spectrum exhibits a prominent band of low frequencies.
If the spectrum were pre-multiplied, it would be forced to zero at zero frequency, and a distinct hump would appear between $St=0.01$ and $St=0.05$. This frequency range is typically associated with breathing when considering pre-multiplied spectra (PoggieAIAA2015; MohammedTaifourJfM2021). In the non–pre-multiplied representation, however, the response exhibits little frequency selectivity for $St\lesssim 0.05$. This is characteristic of a low-pass filter, which is a key feature of low-frequency breathing dynamics (Cura2024; MohammedTaifur_Weiss_2016). This concept traces back to Plotkin1975, who first proposed it in the context of shock oscillations in turbulent SBLI flows. PoggieAIAA2015 later demonstrated its relevance to a larger number of SBLI flows and MohammedTaifur_Weiss_2016; MohammedTaifourJfM2021 extended the concept to a low-speed TSB case which was also observed by Cura2024.
@@ 4.2 Linear stability analysis of the 2D mean field
A linear stability analysis (LSA) is performed on the mean flow in the symmetry plane. The analysis is conducted for spanwise wavenumbers in the range $\beta L_{b}/2\pi=0$22, using the perturbation ansatz given in Eq. 2 and the two-dimensional RANS mean field as base flow, as described in Section 3.1. The eddy-viscosity field from the RANS simulation is used with a scaling factor of $\gamma=3$. As detailed in Appendix A, this scaling was needed to stabilize the base flow to enable resolvent analysis.
Figure 9 shows the computed eigenvalue spectra across the spanwise wavenumber range. The maximum growth rates consistently occur at zero real frequency, $\omega_{r}=0$, corresponding to stationary modes. The wavenumber corresponding to the maximum growth rate is close to $L/\lambda_{z}=1$, indicating that the most amplified spanwise wavelength is approximately equal to the channel width $L$.
@@
The eigenmode with maximum growth rate is shown in Figure 10. It corresponds to a stationary mode ($\omega_{r}=0$) at $\beta\frac{L_{b}}{2\pi}=0.292$ and is positioned above the separation bubble, extending from the separation region to well downstream. The mode shape closely resembles the leading SPOD mode obtained in the symmetry plane (see Fig. 13(a ii and iii)), indicating that the breathing motion is likely associated with a stationary modal mechanism whose growth rate depends strongly on the spanwise wavenumber. These findings are consistent with recent studies by Cura2024; SarrasJfM2024; SavarinoJFM2025.
@@ 4.3 Resolvent analysis of the 2D mean field
Resolvent analysis is performed on the mean field in the symmetry plane. The analysis is conducted for a spanwise wavenumbers range between $\beta\frac{L_{b}}{2\pi}=0$ and $\beta\frac{L_{b}}{2\pi}=2$ and frequency range between $St=0$ and $St=1$. We use the ansatz described in Eq. 6 and the two-dimenstional RANS mean field as base flow, as described in Section 3.1. For each combination of $St$ and $\beta$, the resolvent analysis provides a gain value $\sigma$ and the corresponding forcing and response modes. It is important to note that, since the base flow is assumed to be homogeneous in spanwise direction, this ansatz does not take into account the boundary conditions at the sidewalls.
@@
The low-pass-filter trend of the SPOD eigenvalues is reproduced by the resolvent gain; however, the cutoff frequency is underpredicted. This discrepancy is expected given the following factors. First, the resolvent model predicts low-rank dynamics over a broad range of spanwise wavenumbers, implying that the leading symmetry-plane SPOD eigenvalue spectrum results from contributions of multiple modes and therefore cannot be matched by a single resolvent mode at a single wavenumber; frequency-dependent contributions from several wavenumbers are expected. Second, SPOD and resolvent analyses are theoretically identical only under white-noise forcing (Towne2018). Since the true forcing statistics are unknown, some deviation between the two spectra is expected. Given these limitations, the level of agreement between SPOD eigenvalues and resolvent gain is considered satisfactory.
@@
These findings share similarities with previous studies, particularly with the work of Cura2024. Their results also show maximum amplification at low frequencies and nonzero spanwise wavenumbers, along with a secondary peak at higher frequencies (around $St=1$ to $2\times 10^{-1}$), which they attribute to shear-layer modes. However, an important distinction is that Cura2024 considered a different base flow: a pressure-induced turbulent separation bubble over a flat plate. The spanwise wavenumber differs slightly between the two cases: when normalised by the average bubble length $L_{b}$, the spanwise wavenumber at which they find the maximum resolvent gain is about 20 percent larger then the corresponding value in our setup. Additionally, they identify a weakly damped stationary global eigenmode at approximately one half of the wavenumber associated with the maximum resolvent gain in our case.
@@
While the studies by Cura2024 and Hao2023 show clear amplification at nonzero spanwise wavenumbers, the study by Wu2019 does not exhibit dominant modes with similar spanwise characteristics at comparable frequencies. In the spanwise direction, their computational domain is narrower than the length of a separation bubble, $L_{b}$, whereas in our case, the spanwise wavelength with the highest gain is approximately equal to the bubble length ($\lambda_{z}/L_{b}\approx 1$). This suggests that the phenomena observed in our study could not be resolved within the domain constraints of Wu2019. The results of the resolvent analysis show that there is a linear amplification mechanism intrinsic to the 2D mean flow that occurs within certain ranges of spanwise wavenumbers and low frequencies. This dependence on the spanwise wavenumber aligns with previous observations, such as those by Cura2024. Similar behaviour is reported by SarrasJfM2024, who conducted global linear stability analysis on a RANS mean flow. Their findings reveal stationary three-dimensional global modes becoming unstable only within certain ranges of angles of attack and particular spanwise wavenumbers, a feature consistent with the wavenumber selectivity identified in the present resolvent analysis. It is important to note that the gain derived from the 2D mean flow does not take into account the boundary conditions at the walls.
@@ 4.4 Resolvent-based 3D standing wave model
A comparison between the standing wave model and SPOD results in the symmetry plane reveals that the model accurately captures the shape of the dominant mode. The mode extends from upstream of the separation bubble to well beyond the reattachment region. The 3D visualization (Fig. 13(i)) further illustrates that the mode forms a streamwise-elongated structure. Additionally, the standing wave model successfully captures mode dynamics in the horizontal plane for both the leading mode (compare Figs. 8(b) and 13(b) for $\beta\frac{L_{b}}{2\pi}=1/3$ ($L/\lambda_{z}=1$)) and the subleading mode (compare Figs. 8(c) and 13(c) for $\beta\frac{L_{b}}{2\pi}=1/6$ ($L/\lambda_{z}=0.5$)). In both cases, the model accurately predicts the positions of nodes and antinodes. Further comparison with Fig. 5(c ii) confirms that these modes correspond to the two largest possible standing waves that fit within the channel.
Figure 13: Standing wave model ($u$-component) for $St=0.01$ and $\beta\frac{L_{b}}{2\pi}=1/3$ ($L/\lambda_{z}=1$) in symmetry plane (a) and in horizontal plane (b), for $\beta\frac{L_{b}}{2\pi}=1/6$ ($L/\lambda_{z}=0.5$) in horizontal plane (c), and for $\beta\frac{L_{b}}{2\pi}=1/2$ ($L/\lambda_{z}=1.5$) in horizontal plane (d). The panels (i) show 3D isosurfaces of the real part of the mode. The transparent plane in the panel (a i) visualises the location of the symmetry plane and in (b i) , (c i) and (d i) the location of the horizontal plane. Panels (ii) show the real part of the mode and (iii) show the imaginary part of the mode. For normalization the maximum of the $u$ mode magnitude is used. The dashed lines in (b) , (c) and (d) highlight the position of the centerline. Coordinates are scaled with the spanwise domain length $L$.
@@
Figure 14: Alignment between SPOD mode in horizontal plane and resolvent models. The blue, red and green lines show the alignment between the standing wave model and the leading, subleading, and second subleading, SPOD mode, respectively.
@@
Figure 15: Snapshots of a three dimensional total flow simulation using the mean flow with the scaled standing wave model according to Eq. 12. The blue semitransparent iso-surface shows the separation bubble with $u=0$. The black lines show streamlines anchored upstream and downstream of the separation point. The motion is periodic with a full cycle described by $\omega t=2\pi$. Due to the standing wave nature, the second half of the cycle mirrors the first in reverse order. Panel (a) shows the mode for $\beta\frac{L_{b}}{2\pi}=1/3$ ($L/\lambda_{z}=1$) which is the mode that shows closest similarity with the leading SPOD mode in the experimental data. Panel (b) shows the mode for $\beta\frac{L_{b}}{2\pi}=1/6$ ($L/\lambda_{z}=0.5$) and panel (c) for $\beta\frac{L_{b}}{2\pi}=1/2$ ($L/\lambda_{z}=1.5$) which show closest similarity with the second and third SPOD mode in the experimental data.

### v3

@@ 3.5 Standing Wave Model
In the experimental setup described in Chapter 2.1 the side walls are solid. As with the periodic setup, the wavelength is constrained by the channel width, $L$, and the wavenumbers are discrete. The presence of a solid wall, however, introduces an additional constraint. In solid mechanics and acoustics, wave reflection at boundaries is a well-known phenomenon. A simple demonstration is the classic experiment where a wave pulse traveling along a string fixed at one end reflects upon reaching the boundary. Similarly, water waves hitting solid surfaces exhibit reflection, and in acoustics, echoes in large spaces provide an intuitive example of wave reflection. In compressible numerical simulations, boundary conditions incorporating reflection and refraction properties are standard. However, while our setup is incompressible, we still assume reflection of hydrodynamic waves at the walls.
@@
To derive the standing-wave model, we consider a wave with an arbitrary wavenumber $\beta$ that travels in positive spanwise direction and is reflected at the wall at $z=L$, undergoes a change in orientation that changes the sign of the wavenumber, and propagates to $z=0$, where it is reflected again. In the presence of reflecting boundaries at both ends, a wave is always accompanied by a reflected wave with the same wavelength but inverted orientation. The superposition of two waves of opposite directions with the same wavelength results in the formation of a standing wave. The ansatz for the standing-wave model is then given as:
@@
that is constructed of two resolvent modes with opposite spanwise phase velocities. The same approach could be applied to the eigenmodes. To incorporate spanwise boundary conditions, we enforce slip walls. The slip-wall boundary condition constrains the wall-normal velocity perturbation, $w^{\prime}_{sw}$, to zero at $z=0$ and $z=L$. This holds for all spanwise wavenumbers $\beta$ at $z=0$, but only for $\beta=n\pi/L$ , $n\in\mathbb{Z}$, at $z=L$. The allowed wavenumbers in the standing wave model are thus
@@
In summary, boundary conditions impose fundamental restrictions on the allowable wavelengths. Periodic sidewalls, often employed in numerical studies, permit only integer wavenumbers, with the maximum wavelength constrained by the channel width. By drawing an analogy to acoustics and water waves, we assume the sidewalls exhibit a reflective property. In this framework, resolvent-based waves undergo reflection at the sidewalls, leading to the formation of standing waves. The validity of this model and its agreement with experimental data will be demonstrated in the following section.
@@ 4.1 Empirical description of the TSB low-frequency dynamics
The SPOD spectrum of the streamwise velocity fluctuations measured by PIV on the symmetry plane is presented in Figure 6(a) which shows the mode energies of all SPOD modes $\lambda_{i}$. Figure 7(a) shows the energy share of each mode at every frequency which is calculated by dividing the mode energy by the sum of all modes at the given frequency. It reveals a leading mode that has a significantly higher energy share compared to the other modes. This is also known as high-gain separation and describes low-rank dynamics, meaning that the dynamics of the flow are dominated by one or very few coherent structures in the corresponding frequency range (SchmidtColonius2020). This mode concentrates the majority of its energy within the low-frequency range, specifically below $St=0.05$. The spectral resolution of the SPOD is determined by the sampling frequency and the number of snapshots per block. In the symmetry plane, a window size of 200 snapshots yields a minimum resolvable frequency of $St=0.01$. To resolve lower frequencies, we increased the window length to 2000 samples, extending the resolution down to $St=0.001$ (not shown). While this confirmed the presence of the same low-pass filter behavior, the resulting spectra exhibited reduced statistical convergence due to the lower number of available blocks for averaging. A similar low-rank behaviour for low frequencies has been observed by WeissAIAAJ2022 using SPOD of wall-shear stress signals. Instead of a single dominant characteristic frequency, the spectrum exhibits a prominent band of low frequencies.
If the spectrum were pre-multiplied, it would be forced to zero at zero frequency, and a distinct hump would appear between $St=0.01$ and $St=0.05$. This frequency range is typically associated with breathing when considering pre-multiplied spectra (PoggieAIAA2015; MohammedTaifourJfM2021). In the non–pre-multiplied representation, however, the response exhibits little frequency selectivity for $St\lesssim 0.05$. This is characteristic of a low-pass filter, which is a key feature of low-frequency breathing dynamics (Cura2024; MohammedTaifourJfM2021). This concept traces back to Plotkin1975, who first proposed it in the context of shock oscillations in turbulent SBLI flows. PoggieAIAA2015 later demonstrated its relevance to a larger number of SBLI flows and MohammedTaifur_Weiss_2016; MohammedTaifourJfM2021 extended the concept to a low-speed TSB case which was also observed by Cura2024.
@@ 4.2 Linear stability analysis of the 2D mean field
A linear stability analysis (LSA) is performed on the mean flow in the symmetry plane. The analysis is conducted for spanwise wavenumbers in the range $\beta L_{b}/2\pi=0$22, using the perturbation ansatz given in Eq. 2 and the two-dimensional RANS mean field as base flow, as described in Section 3.1.
The eddy viscosity is prescribed as
$\nu_{t}=\gamma\,\nu_{t,\text{RANS}}(x,y),$ (10)
where $\nu_{t,\text{RANS}}$ denotes the eddy viscosity obtained from the RANS solution, and $\gamma$ is a scaling factor. Using the unscaled RANS eddy viscosity ($\gamma=1$) yields unstable eigenvalues. Performing a resolvent analysis on this mean flow would therefore require the discounted resolvent formulation (Rolandi2024_invitation), which we aimed to avoid.
Instead, we conducted a parametric study with different scaling factors to identify a suitable value of $\gamma$. Throughout the study, the maximum growth rates consistently occur at zero real frequency, $\omega_{r}=0$, corresponding to stationary modes. Figure 9(a) shows the stationary eigenvalues ($\omega_{r}=0$) obtained for varying $\gamma$. Increasing $\gamma$ reduces the growth rates of the stationary eigenmodes; for $\gamma\geq 2$, all eigenvalues are stable. The choice of $\gamma$ can be further guided by considering its influence on the frequency response. Bugeat2022JfM and SarrasPhD_thesis demonstrate that the growth rate of a stationary eigenmode determines the cutoff frequency of the associated low-pass filter in the resolvent response. This trend is evident in Fig. 9(b), which shows the resolvent gain for $\beta L_{b}/2\pi=1/3$ ($L/\lambda_{z}=1$), corresponding to the most energetic structure in the SPOD. As $\gamma$ increases beyond 2, the low-frequency plateau of the gain curve broadens, and the roll-off shifts to higher frequencies. This behaviour is characteristic of an increased low-pass filter cutoff frequency. For the present analysis, $\gamma=3$ is selected as an appropriate compromise, providing a realistic cutoff frequency of $St\approx 0.01$ while avoiding the numerical difficulties associated with excessively high eddy-viscosity values. Note that a more rigorous approach to avoid the tuning of eddy viscosity would involve assimilating the RANS solution to the measured mean field and incorporating a linearized eddy-viscosity model into the linear operator (SarrasJfM2024). This, however, is beyond the scope of the present study. The close agreement between SPOD and linearized results justifies our approach.
Figure 9: (a) Stationary eigenvalues ($St_{r}=0$) computed for five eddy-viscosity scaling factors, $\nu_{t}(x,y)=\gamma\,\nu_{t,\text{RANS}}(x,y)$ with $\gamma\in[1,2,3,4,5]$. (b) Leading resolvent gains for $\beta L_{b}/2\pi=1/3$ (equivalently $L/\lambda_{z}=1$) compared with the SPOD spectrum in the symmetry plane for the same set of $\gamma$ values.
Figure 10 shows the computed eigenvalue spectra for $\gamma=3$ across the spanwise wavenumber range. The maximum growth rates consistently occur at zero real frequency, $\omega_{r}=0$, corresponding to stationary modes. The wavenumber corresponding to the maximum growth rate is close to $L/\lambda_{z}=1$, indicating that the most amplified spanwise wavelength is approximately equal to the channel width $L$.
@@
The eigenmode with maximum growth rate is shown in Figure 11. It corresponds to a stationary mode ($\omega_{r}=0$) at $\beta\frac{L_{b}}{2\pi}=0.292$ and is positioned above the separation bubble, extending from the separation region to well downstream. The mode shape closely resembles the leading SPOD mode obtained in the symmetry plane (see Fig. 14(a ii and iii)), indicating that the breathing motion is likely associated with a stationary modal mechanism whose growth rate depends strongly on the spanwise wavenumber. These findings are consistent with recent studies by Cura2024; Cura2025; SavarinoJFM2025.
@@ 4.3 Resolvent analysis of the 2D mean field
Resolvent analysis is performed on the mean field in the symmetry plane. The analysis is conducted for a spanwise wavenumbers range between $\beta\frac{L_{b}}{2\pi}=0$ and $\beta\frac{L_{b}}{2\pi}=2$ and frequency range between $St=0$ and $St=1$. We use the ansatz described in Eq. 6, the two-dimensional RANS mean field as base flow, as described in Section 3.1, and an eddy viscosity scaling factor $\gamma=3$, as described in Section 4.2. For each combination of $St$ and $\beta$, the resolvent analysis provides a gain value $\sigma$ and the corresponding forcing and response modes. It is important to note that, since the base flow is assumed to be homogeneous in spanwise direction, this ansatz does not take into account the boundary conditions at the sidewalls.
@@
The low-pass-filter trend of the SPOD eigenvalues is reproduced by the resolvent gain; however, the cutoff frequency is underpredicted. This discrepancy is expected given the following factors. First, the resolvent model predicts low-rank dynamics over a broad range of spanwise wavenumbers, implying that the leading symmetry-plane SPOD eigenvalue spectrum results from contributions of multiple modes and therefore cannot be matched by a single resolvent mode at a single wavenumber; frequency-dependent contributions from several wavenumbers are expected. Second, SPOD and resolvent analyses are theoretically identical only under white-noise forcing (Towne2018). Since the true forcing statistics are unknown, some deviation between the two spectra is expected. Third, as shown in Fig. 9(b), the cutoff frequency is influenced by the eddy viscosity scaling $\gamma$. Given these limitations, the level of agreement between SPOD eigenvalues and resolvent gain is considered satisfactory.
@@
The findings presented in Fig. 12 share similarities with previous studies, particularly with the work of Cura2024. Their results also show maximum amplification at low frequencies and nonzero spanwise wavenumbers, along with a secondary peak at higher frequencies (around $St=1$ to $2\times 10^{-1}$), which they attribute to shear-layer modes. However, an important distinction is that Cura2024 considered a different base flow: a pressure-induced turbulent separation bubble over a flat plate. The spanwise wavenumber differs slightly between the two cases: when normalised by the average bubble length $L_{b}$, the spanwise wavenumber at which they find the maximum resolvent gain is about 20 percent larger than the corresponding value in our setup. Additionally, they identify a weakly damped stationary global eigenmode at approximately one half of the wavenumber associated with the maximum resolvent gain in our case.
@@
While the studies by Cura2024 and Hao2023 show clear amplification at nonzero spanwise wavenumbers, the study by Wu2019 does not exhibit dominant modes with similar spanwise characteristics at comparable frequencies. In the spanwise direction, their computational domain is narrower than the length of a separation bubble, $L_{b}$, whereas in our case, the spanwise wavelength with the highest gain is approximately equal to the bubble length ($\lambda_{z}/L_{b}\approx 1$). This suggests that the phenomena observed in our study could not be resolved within the domain constraints of Wu2019. The results of the resolvent analysis show that there is a linear amplification mechanism intrinsic to the 2D mean flow that occurs within certain ranges of spanwise wavenumbers and low frequencies. This dependence on the spanwise wavenumber aligns with previous observations, such as those by Cura2024. Similar behaviour is reported by SarrasJfM2024, who conducted global linear stability analysis on a RANS mean flow. Their findings reveal stationary three-dimensional global modes becoming unstable only within certain ranges of angles of attack and particular spanwise wavenumbers, a feature consistent with the wavenumber selectivity identified in the present resolvent analysis. It is important to note that the gain derived from the 2D mean flow does not take into account the boundary conditions at the walls. Recent results by Cura2025 indicate that variations in reverse flow and separation-bubble length modify the growth rates, whereas the dominant spanwise wavenumbers are much less sensitive. This may explain why experimental and modelling results show good agreement despite differences in reverse flow and separation-bubble length, as discussed in Section 3.1.
@@ 4.4 Resolvent-based 3D standing wave model
A comparison between the standing wave model and SPOD results in the symmetry plane reveals that the model accurately captures the shape of the dominant mode. The mode extends from upstream of the separation bubble to well beyond the reattachment region. The 3D visualization (Fig. 14(i)) further illustrates that the mode forms a streamwise-elongated structure. Additionally, the standing wave model successfully captures mode dynamics in the horizontal plane for both the leading mode (compare Figs. 8(b) and 14(b) for $\beta\frac{L_{b}}{2\pi}=1/3$ ($L/\lambda_{z}=1$)) and the subleading mode (compare Figs. 8(c) and 14(c) for $\beta\frac{L_{b}}{2\pi}=1/6$ ($L/\lambda_{z}=0.5$)). In both cases, the model accurately predicts the positions of nodes and antinodes. Further comparison with Fig. 5(c ii) confirms that these modes correspond to the two largest possible standing waves that fit within the channel.
Figure 14: Standing wave model ($u$-component) for $St=0.01$ and $\beta\frac{L_{b}}{2\pi}=1/3$ ($L/\lambda_{z}=1$) in symmetry plane (a) and in horizontal plane (b), for $\beta\frac{L_{b}}{2\pi}=1/6$ ($L/\lambda_{z}=0.5$) in horizontal plane (c), and for $\beta\frac{L_{b}}{2\pi}=1/2$ ($L/\lambda_{z}=1.5$) in horizontal plane (d). The panels (i) show 3D isosurfaces of the real part of the mode. The transparent plane in the panel (a i) visualises the location of the symmetry plane and in (b i), (c i), and (d i) the location of the horizontal plane. Panels (ii) show the real part of the mode and (iii) show the imaginary part of the mode. For normalization the maximum of the $u$ mode magnitude is used. The dashed lines in (b), (c), and (d) highlight the position of the centerline. Coordinates are scaled with the spanwise domain length $L$.
@@
Figure 15: Alignment between SPOD mode in horizontal plane and resolvent models for St=St=0.01. The blue, red and green lines show the alignment between the standing wave model and the leading, subleading, and second subleading, SPOD mode, respectively.
@@
Figure 16: Snapshots of a three dimensional total flow simulation for $St=$ 0.01 using the mean flow with the scaled standing wave model according to Eq. 13. The blue semitransparent iso-surface shows the separation bubble with $u=0$. The black lines show streamlines anchored upstream and downstream of the separation point. The motion is periodic with a full cycle described by $\omega t=2\pi$. Due to the standing wave nature, the second half of the cycle mirrors the first in reverse order. Panel (a) shows the mode for $\beta\frac{L_{b}}{2\pi}=1/3$ ($L/\lambda_{z}=1$) which is the mode that shows closest similarity with the leading SPOD mode in the experimental data. Panel (b) shows the mode for $\beta\frac{L_{b}}{2\pi}=1/6$ ($L/\lambda_{z}=0.5$) and panel (c) for $\beta\frac{L_{b}}{2\pi}=1/2$ ($L/\lambda_{z}=1.5$) which show closest similarity with the second and third SPOD mode in the experimental data.